Beds Price per Bed Average Income Price of Product A Price of Product B
200 1,000 50,000 400 400
300 750 60,000 500 300

1. calculate the arc price elasticity for beds
2. calculate the arc income elasticity for beds
1
Expert's answer
2020-11-12T17:27:06-0500

1. The arc price elasticity for beds is:

"Ed = \frac{300 - 200} {750 - 1,000} × \frac{750 + 1,000} {300 + 200} = - 1,750/1,250 = -1.4."

2. The arc income elasticity for beds is:

"Ed = \frac{60,000 - 50,000} {750 - 1,000} × \frac{750 + 1,000} {60,000 + 50,000} = -7/11 = -0.64."